OpenStudy (anonymous):

How would I put (-144) to the 1/2 power in radical form?

OpenStudy (anonymous):

this is #13

OpenStudy (anonymous):

1/2 power is square root.

OpenStudy (anonymous):

Can you explain to me how you do this step by step?

OpenStudy (anonymous):

i^2=-1(-144)^1/2=(12*12 i^2)^1/2=12i

OpenStudy (anonymous):

Why is there a variable in the equation?

OpenStudy (anonymous):

i=for use -ve

OpenStudy (anonymous):

i=imag

OpenStudy (anonymous):

I'm sorry but I still don't understand. Can you explain this to me?

OpenStudy (anonymous):

your ans

OpenStudy (anonymous):

\(\large (-144)^{1/2} = \sqrt{-144}\)

OpenStudy (anonymous):

sister your book ans

OpenStudy (anonymous):

\(\large \sqrt{-144}=\sqrt{-1} \cdot \sqrt{144}\) \(\large \sqrt{-1} = i\), \(\large \sqrt{144}=12.\)

OpenStudy (anonymous):

oh okay, thank you. can you guys explain to me how to simplify the expression 16^3/4 if possible?

OpenStudy (anonymous):

Rational exponents are 'power-over-root' \(\large 16^{3/4} = \sqrt[4]{(16)^3}\)

OpenStudy (anonymous):

16^(3/4)=(2^4)^(\[16^(3/4)=(2^4)^(3/4)=2^((4\times3)/4)=2^3=8

OpenStudy (anonymous):

^wow, what a way to over-complicate something that is really simple. :-p

OpenStudy (anonymous):

\(\large 16^{3/4}=(\sqrt[4]{16})^3\) \(\large \sqrt[4]{16} =2\) \(\large 2^3=8\)

OpenStudy (anonymous):

16^1-1/4=16^1/16^1/4=16/2=8

OpenStudy (anonymous):

\[(-144)^{\frac{1}{2}}=12 \sqrt{-1}=12 i \]\[-144^{\frac{1}{2}}=-12 \]